OpenStudy (anonymous):

Multiply and simplify A). (b+2)(b+2) B). (x+3)(x-3)

OpenStudy (anonymous):

a) b^2+2b+2b+4 -> b^2+4b+4 b) x^2-3x+3x-9 -> x^2-9

OpenStudy (anonymous):

Thank you so much.!:)

OpenStudy (anonymous):

(b+2)(b+2) = (b+2)² = b² +2(b)(2) + 2² = b² + 4b + 4 RULES --> (a+b)²=a² + 2ab + b² AND (a-b)²=a² - 2ab + b² just like (x+3)(x+3)=(x+3)² = x² + 2(x)(3) + 3² = x² + 6x + 9 and (c-5)(c-5)=(c-5)² = c² - 2(c)(5) + (5)² = c² - 10c + 25 (x+3)(x-3) = x² - 3²=x² - 9. The rule is below... (a-b)(a+b)=a² - b²